OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x – 2y = 2 x + 3y = 17

OpenStudy (luffingsails):

So, solve the first equation for x: x = 2 + 2y Now, substitute that equation into the second equation for x... x + 3y = 17 (2+2y) + 3y = 17 2 + 5y = 17 5y = 17 - 2 = 15 y = 3 Substitute that back into any of our equations to solve for x. I'll use x = 2 + 2y x = 2 + 2(3) = 8. So, x = 8, y = 3 is our solution. Make sense?

OpenStudy (across):

You can also solve this problem using linear algebra \[\begin{bmatrix} 1 & -2\\ 1 & 3 \end{bmatrix}\begin{bmatrix} x\\ y \end{bmatrix}=\begin{bmatrix} 2\\ 17 \end{bmatrix}\]Matrix determinant \[det=\frac{1}{5}\]Matrix minor \[\begin{bmatrix} 3 & 1\\ -2 & 1 \end{bmatrix}\]Matrix cofactor \[\begin{bmatrix} 3 & -1\\ 2 & 1 \end{bmatrix}\]Matrix transpose \[\begin{bmatrix} 3 & 2\\ -1 & 1 \end{bmatrix}\]Putting it all together \[\begin{bmatrix} x\\ y \end{bmatrix}=\frac{1}{5}\begin{bmatrix} 3 & 2\\ -1 & 1 \end{bmatrix}\begin{bmatrix} 2\\ 17 \end{bmatrix}=\frac{1}{5}\begin{bmatrix} 40\\ 15 \end{bmatrix}=\begin{bmatrix} 8\\ 3 \end{bmatrix}\]Therefore, \[x=8, y=3\]Sorry, I'm bored waiting for class. lol